# Search for a diffuse flux of photons with energies above tens of PeV at the Pierre Auger Observatory

### v2

@@ 4 Discrimination between photon and proton events
The search for primary photons in the cosmic-ray flux is a classification problem of a tiny signal in the overwhelmingly dominant hadronic background. The muon density is a powerful observable for distinguishing between air showers dominated by the electromagnetic component and those with significant hadronic contributions. Using the simulated set described in Section 3.1, we define in Section 4.1 a muon-based observable that can be related to the mass of the primary cosmic ray. We optimize it to maximize the separation power between photon- and proton-initiated events in Section 4.2, where the energy is assigned with the photon-equivalent energy scale, as discussed in Section 3.4.
@@ 4.1 Muon content estimator, $M_{b}$
Each measured muon density is normalized by the average muon density, $\rho_{\text{pr}}$, expected in proton-initiated events at a predefined reference distance, $r_{\text{pr}}$, and is parametrized in terms of the primary energy and zenith angle. This normalization serves a dual purpose in the discrimination method. First, it offers an initial estimate of the muon content, with hadronic-induced showers typically yielding positive values of $M_{b}$ and photon-initiated showers producing negative values. Second, by normalizing each muon density by $\rho_{\text{pr}}$, the energy and angular dependencies of $M_{b}$ are largely absorbed, while the composition-dependent scaling of the muon densities is preserved. The power-law index, $b$, is selected to maximize the separation power of $M_{b}$ between photonic and hadronic events, as discussed in Section 4.2.
The reference distance, $r_{\text{pr}}$, is chosen to minimize fluctuations in the number of muons intersecting the sensitive area of a UMD station. The average number of muons, $\overline{N_{\mu}}$, decreases with the distance to the shower axis, following the lateral spread of the shower front. Conversely, the relative fluctuations, $\sigma_{\mu}/\overline{N_{\mu}}$, increase with distance, as shown in Fig. 7, left. These relative fluctuations are influenced by both the shower-to-shower and the Poissonian fluctuations sprouting from the decreasing number of impinging muons [66]. The relative fluctuations are estimated as the root-mean-square of the $N_{\mu}$ distributions in distance bins [^3] At distances below $\sim$300 m, shower-to-shower fluctuations dominate in the energy range under study, exceeding the Poissonian fluctuations by a factor between $2$ and $3$ at distances between $10^{16.5}$ eV and $10^{17}$ eV.. A plateau in the relative fluctuations is observed up to $\sim 300$ m, mildly dependent on the primary energy and zenith angle. Consequently, we select a reference distance of $r_{\text{pr}}=200\,\mathrm{m}$, which corresponds to about $25\%$ statistical fluctuations in the number of muons with a minimal variation across the studied energy range.
The normalization factor, $\rho_{\text{pr}}$, is then defined as the average of the muon density in proton-induced events at 200 m from the shower axis. It is parameterized in terms of the proton energy, $E_{\text{pr}}$, and zenith angle using the muon densities measured between $195$ m and $205$ m from the shower axis. The energy dependence of $\rho_{\text{pr}}$ is illustrated in Fig. 7, right, and can be described by:
@@
The estimated values of the free parameters, $\rho_{0}$ and $c$, using events in different zenith angle intervals are depicted in Fig. 8. The normalization, $\rho_{0}$, decreases with the zenith angle, representing the atmospheric attenuation of the muon component. In contrast, the power-law index, $c$, governs the energy-driven muon production and shows no significant angular dependence. A quadratic dependence on $\cos^{2}\theta-\cos^{2}30\,\mathrm{{}^{\circ}}$ is proposed to describe $\lg(\rho_{0})$, while $c$ is assumed to be a constant. These four free parameters are estimated using a maximum likelihood approach to the unbinned data and presented in Table 3.
@@
Figure 8: The free parameters of Eq. 4.2 estimated by $\chi^{2}$ minimization to the muon densities in angular bins with superimposed fitted quadratic and constant models.
@@
Mean $-0.108$ $0.262$ $-0.591$ $0.890$
[…]
Table 3: The constant, linear and quadratic coefficients in $\cos^{2}\theta-\cos^{2}30\,\mathrm{{}^{\circ}}$ describing $\lg(\rho_{0}(\theta))$ in Eq. 4.2. The last column corresponds to the parameter $c$ present in Eq. 4.2.
@@ 4.2 Optimization of $M_{b}$
Because the number of muons on the ground decreases with increasing distance from the shower axis, the observed muon densities exhibit significant statistical uncertainties at large distances due to the finite detector size, independent of the primary particle. Therefore, $M_{b}$ is defined using the measurements taken with the “hottest” station (the one with the highest SD signal) and its six nearest neighbors, which together define the “hottest” hexagon of an event. An example of the $M_{b=1}$ distributions for photon- and proton-initiated events is displayed in Fig. 9, left. While there is a clear separation between the two populations, a subset of muon-poor proton events is visible. These events generally correspond to showers with a dominant electromagnetic component, generated by the decay of leading $\pi^{0}$ produced in first interactions and carrying a substantial fraction of the primary cosmic-ray energy. Vice versa, a minor subset of photon showers with hadronic-like muon content is also observed. These events result from the decay of a leading $\pi^{\pm}$, caused by a photonuclear interaction, which initiates a hadronic sub-shower early in the shower development.
@@
The choice of $b$ in Eq. 4.1 aims to minimize the risk of misclassifying a background event as a photon event, while still ensuring a substantial probability of identifying photon events. The hadronic background is assumed purely of protons to provide a conservative estimation of the photon-hadron separation power, given the existence of a non-negligible contribution of primaries heavier than protons in the cosmic-ray flux [12, 14]. We define background contamination as the ratio of the number of proton events below a specified $M_{b}$ threshold to the total number of proton events. To estimate the former from the available statistics, the $M_{b}$ of the $10\%$ most photon-like proton events are fitted using an unbinned maximum likelihood procedure, assuming an exponential model (shown by the black line in Fig. 9, left). Then the number of proton events below a given $M_{b}$ threshold is calculated from the integral of the fitted tail [^4] This estimation of the background contamination does not depend on the size of the most photon-like background tail, i.e., it remains stable when considering between $5\%$ and $15\%$ of the proton events in the definition of the background tail, thus adding to the robustness of the method.. The signal efficiency, on the other hand, is defined as the ratio between the photon-initiated events having values of $M_{b}$ smaller than the threshold value and the total number of photon events.
@@
As shown in Fig. 9, right, the contamination is minimal for a range of values of $b$ around unity in the energy range of interest for a fixed signal efficiency of $50\%$. It decreases with increasing energy due to the enhanced air-shower muon content. Therefore, $b=1$ is chosen as the index that minimizes the background contamination, defining the discrimination observable as $M_{1}$.
@@
The background contamination and signal efficiency can be computed by scanning over all possible values of the $M_{1}$ threshold, as exemplified in Fig. 10. The contamination decreases with increasing primary energy, reflecting the increasing number of muons produced in hadronic air showers at higher energies. In addition, a larger signal efficiency can be reached at the expense of a larger background contamination.
@@
We note that the background contamination and signal efficiency shown in Fig. 10 are obtained assuming a complete UMD hexagon, e.g., the north-west hexagon in Fig. 1, left. We verified that the choice of $b=1$ remains effective even when considering missing UMD stations in the first ring of neighbors. However, the discrimination power, i.e. both the background contamination and signal efficiency, varies depending on the real configurations of the detector, as discussed in the next section.
@@ 5 Selection of photon events in data
In this section, we study the application of the discrimination method to data, considering the actual configuration of the detector. Given that the fraction of photon primaries expected to be present in the cosmic-ray flux may be well below $10^{-5}$, the standard event reconstruction, fine-tuned to hadronic events, is employed for the data. The event selection follows the procedure described in Section 2. Both the estimated shower size and zenith angle are used to assign photon-equivalent energy to each event.
The observable $M_{1}$ quantifies the muon content of each event, as defined in Section 4.2, and scales with the number of available UMD stations in the hottest hexagon. Air showers with low muon content, charactistic of photon-like events, would have smaller values of $M_{1}$ compared to hadronic events. However, if any UMD stations are missing, either due to not being deployed yet at the time of detection or to technical issues during data acquisition, less information on the muon content would be available. This results in smaller values of $M_{1}$ when fewer than seven UMD stations are operational, leading to an increase of the likelihood of misidentifying background events as photon-initiated. Thus, before applying the discrimination method to the selected data set, it is important to evaluate the impact of missing UMD modules across the seven stations of a hexagon.
Due to the irregular placement of UMD stations in the SD-433, as seen in Fig. 1, left, the muon component arriving at the ground is sampled by a variable number of UMD stations. Therefore, the threshold value of $M_{1}$ to classify an event as a photon candidate, i.e., the photon candidate cut $M_{1}^{\text{cut}}$, is obtained according to the possible configurations in the hottest hexagon. The events are classified in six categories, as listed in Table 4, depending on the following three criteria: the number of available UMD stations (i.e., stations with at least one active module), the total area spanned by these stations, and the relative position of any missing stations (i.e., stations without active modules). The hottest station must contain three active UMD modules regardless of the event category.
Cat. No. of first-ring UMD stations Minimum UMD area in first ring (m2) Missing UMD stations in first ring No. of events
I $6$ $190$ 00 3,295
II $5$ $140$ $1$ 1,491
III $4$ $110$ $2$ non-NN 8,417
IV $4$ $110$ $2$ NN 298
V $3$ $80$ $3$ non-NN 1,016
VI $3$ $80$ $2$ NN + $1$ non-NN 1,402
@@
The discrimination power of a hexagonal grid that mimics the SD-433 is evaluated through simulations (see Section 3.1), assuming each SD station is paired with a 30 m2 UMD station, except for the aforementioned 50 m2 station. UMD stations or individual modules are randomly masked to simulate missing measurements and assess the discrimination power for each event category. $M_{1}^{\text{cut}}$ is defined as the threshold value at which the signal efficiency reaches $50\%$, determined as the median of the $M_{1}$ distribution for simulated photon events in energy bins. Such signal efficiency is suitable to reach background contamination levels smaller than $10^{-4}$ in all event categories. Therefore, an event is tagged as a photon candidate if $M_{1}<M_{1}^{\text{cut}}$. As shown in Fig. 11, left, $M_{1}^{\text{cut}}$ increases with primary energy and with the number of UMD stations in the hottest hexagon . The evolution of $M_{1}^{\text{cut}}$ is described by a function of the logarithmic energy for each event category and reported in Appendix B. As depicted in Fig. 11, right, the contamination remains below $10^{-5}$ for events in category I, while it increases as the number of first-ring UMD stations decreases, with the relative position of missing UMD stations having a significant impact. For instance, the contamination in events with two missing neighboring stations is two to four times larger than in events with non-neighboring missing stations. The background contamination is modeled using an exponential function of logarithmic energy, as detailed in Appendix B.
@@
Figure 11: The photon candidate cut (left) and the background contamination at $50\%$ signal efficiency (right) in terms of the photon-equivalent energy for each of the event categories cited in Table 4. The dashed lines show examples of the fitted power-law and exponential models of the logarithmic energy, respectively.
@@
The parameterization of the background contamination for each event category can be employed to obtain a conservative estimation of the number of hadronic events in data that would be misidentified as photon candidates. For instance, the expected number of fake photon events in the category I subset can be computed as $(4.04\pm 0.04)\times 10^{-3}$ by integrating the blue curve in Fig. 11, right, above $10^{16.7}$ eV [68]. Since 3,295 events above $10^{16.7}$ eV fall in that event category, this is equivalent to one fake photon event approximately every 815,650 events [^5] The expected number of fake photon events in the category IV and VI subsets, corresponding to the configurations with the largest background contamination, is $3.12\times 10^{-3}$ and $8.20\times 10^{-3}$ respectively. It can be translated to one fake photon event every 95,460 and 170,890 events in each of the subsets.. Considering that the category I events were acquired over $\sim 10\,\mathrm{months}$ of the observation time, this suggests that one fake photon event would be expected every 106 years per hexagon.
@@
Given the low probability of observing a background event consistent with the photon hypothesis, a fraction of the selected data, referred to as the burnt data sample, is unblinded to check that the $M_{1}$ distribution in data agrees with simulations, before carrying on the unblinding of the full selected data set. The burnt data represent a fraction $f_{\text{burnt}}=0.1$ of the selected data set. The $M_{1}$ distributions of burnt data for each event category are compared with the simulated distributions, as exemplified in Fig. 12. Simulated events are weighted by $E_{\gamma,\text{eq}}^{-2}$ to obtain energy-integrated distributions, while no weighting is applied to the burnt data sample. As expected, the $M_{1}$ distribution for the burnt data is compatible with a hadronic origin. It is important to note that the photon-equivalent energy assigned to a data event is around $\sim 30\%$ higher than that assigned using the hadronic-optimized energy scale. This discrepancy arises from the larger muon content in cosmic-ray showers compared to the photon showers used to define the photon-equivalent energy calibration. As a result, the reference muon density $\rho_{\text{pr}}$ used in the $M_{1}$ definition (see Eq. 4.1) is shifted to higher values in data compared to simulations, leading to smaller $M_{1}$ values in the former case. Thus, the distribution for data is artificially shifted towards the simulated proton distribution. Nevertheless, the $M_{1}$ observable is optimized to discriminate between events with significantly different muonic components -such as photons versus all hadronic species- due to the logarithm scaling in its definition, making it largely insensitive to the primary hadronic mass. This limitation does not impact the search for photon events in data, as discussed in the next section.

### v3

@@ 4 Discrimination between photon and proton events
The search for primary photons in the cosmic-ray flux is a classification problem of a tiny signal in the overwhelmingly dominant hadronic background. The muon density is a powerful observable for distinguishing between air showers dominated by the electromagnetic component and those with significant hadronic contributions. Using the simulated set described in Section˜3.1, we define in Section˜4.1 a muon-based observable that can be related to the mass of the primary cosmic ray. We optimize it to maximize the separation power between photon- and proton-initiated events in Section˜4.2, where the energy is assigned with the photon-equivalent energy scale, as discussed in Section˜3.4.
@@ 4.1 Muon content estimator, $M_{b}$
Each measured muon density is normalized by the average muon density, $\rho_{\text{pr}}$, expected in proton-initiated events at a predefined reference distance, $r_{\text{pr}}$, and is parametrized in terms of the primary energy and zenith angle. This normalization serves a dual purpose in the discrimination method. First, it offers an initial estimate of the muon content, with hadronic-induced showers typically yielding positive values of $M_{b}$ and photon-initiated showers producing negative values. Second, by normalizing each muon density by $\rho_{\text{pr}}$, the energy and angular dependencies of $M_{b}$ are largely absorbed, while the composition-dependent scaling of the muon densities is preserved. The power-law index, $b$, is selected to maximize the separation power of $M_{b}$ between photonic and hadronic events, as discussed in Section˜4.2.
The reference distance, $r_{\text{pr}}$, is chosen to minimize fluctuations in the number of muons intersecting the sensitive area of a UMD station. The average number of muons, $\overline{N_{\mu}}$, decreases with the distance to the shower axis, following the lateral spread of the shower front. Conversely, the relative fluctuations, $\sigma_{\mu}/\overline{N_{\mu}}$, increase with distance, as shown in Fig.˜7, left. These relative fluctuations are influenced by both the shower-to-shower and the Poissonian fluctuations sprouting from the decreasing number of impinging muons [66]. The relative fluctuations are estimated as the root-mean-square of the $N_{\mu}$ distributions in distance bins [^3] At distances below $\sim$300 m, shower-to-shower fluctuations dominate in the energy range under study, exceeding the Poissonian fluctuations by a factor between $2$ and $3$ at distances between $10^{16.5}$ eV and $10^{17}$ eV.. A plateau in the relative fluctuations is observed up to $\sim 300$ m, mildly dependent on the primary energy and zenith angle. Consequently, we select a reference distance of $r_{\text{pr}}=200\,\mathrm{m}$, which corresponds to about $25\%$ statistical fluctuations in the number of muons with a minimal variation across the studied energy range.
The normalization factor, $\rho_{\text{pr}}$, is then defined as the average of the muon density in proton-induced events at 200 m from the shower axis. It is parameterized in terms of the proton energy, $E_{\text{pr}}$, and zenith angle using the muon densities measured between $195$ m and $205$ m from the shower axis. The energy dependence of $\rho_{\text{pr}}$ is illustrated in Fig.˜7, right, and can be described by:
@@
The estimated values of the free parameters, $\rho_{0}$ and $c$, using events in different zenith angle intervals are depicted in Fig.˜8. The normalization, $\rho_{0}$, decreases with the zenith angle, representing the atmospheric attenuation of the muon component. In contrast, the power-law index, $c$, governs the energy-driven muon production and shows no significant angular dependence. A quadratic dependence on $\cos^{2}\theta-\cos^{2}30\,\mathrm{{}^{\circ}}$ is proposed to describe $\lg(\rho_{0})$, while $c$ is assumed to be a constant. These four free parameters are estimated using a maximum likelihood approach to the unbinned data and presented in Table˜3.
@@
Figure 8: The free parameters of Eq.˜4.2 estimated by $\chi^{2}$ minimization to the muon densities in angular bins with superimposed fitted quadratic and constant models.
@@
Mean 0.108-0.108 $0.262$ 0.591-0.591 $0.890$
Uncertainty $\phantom{+}0.002$ $\phantom{+}0.010$ $\phantom{+}0.055$ $\phantom{+}0.006$
Table 3: The constant, linear and quadratic coefficients in $\cos^{2}\theta-\cos^{2}30\,\mathrm{{}^{\circ}}$ describing $\lg(\rho_{0}(\theta))$ in Eq.˜4.2. The last column corresponds to the parameter $c$ present in Eq.˜4.2.
@@ 4.2 Optimization of $M_{b}$
Because the number of muons on the ground decreases with increasing distance from the shower axis, the observed muon densities exhibit significant statistical uncertainties at large distances due to the finite detector size, independent of the primary particle. Therefore, $M_{b}$ is defined using the measurements taken with the “hottest” station (the one with the highest SD signal) and its six nearest neighbors, which together define the “hottest” hexagon of an event. An example of the $M_{b=1}$ distributions for photon- and proton-initiated events is displayed in Fig.˜9, left. While there is a clear separation between the two populations, a subset of muon-poor proton events is visible. These events generally correspond to showers with a dominant electromagnetic component, generated by the decay of leading $\pi^{0}$ produced in first interactions and carrying a substantial fraction of the primary cosmic-ray energy. Vice versa, a minor subset of photon showers with hadronic-like muon content is also observed. These events result from the decay of a leading $\pi^{\pm}$, caused by a photonuclear interaction, which initiates a hadronic sub-shower early in the shower development.
@@
The choice of $b$ in Eq.˜4.1 aims to minimize the risk of misclassifying a background event as a photon event, while still ensuring a substantial probability of identifying photon events. The hadronic background is assumed purely of protons to provide a conservative estimation of the photon-hadron separation power, given the existence of a non-negligible contribution of primaries heavier than protons in the cosmic-ray flux [12, 14]. We define background contamination as the ratio of the number of proton events below a specified $M_{b}$ threshold to the total number of proton events. To estimate the former from the available statistics, the $M_{b}$ of the $10\%$ most photon-like proton events are fitted using an unbinned maximum likelihood procedure, assuming an exponential model (shown by the black line in Fig.˜9, left). Then the number of proton events below a given $M_{b}$ threshold is calculated from the integral of the fitted tail [^4] This estimation of the background contamination does not depend on the size of the most photon-like background tail, i.e., it remains stable when considering between $5\%$ and $15\%$ of the proton events in the definition of the background tail, thus adding to the robustness of the method.. The signal efficiency, on the other hand, is defined as the ratio between the photon-initiated events having values of $M_{b}$ smaller than the threshold value and the total number of photon events.
@@
As shown in Fig.˜9, right, the contamination is minimal for a range of values of $b$ around unity in the energy range of interest for a fixed signal efficiency of $50\%$. It decreases with increasing energy due to the enhanced air-shower muon content. Therefore, $b=1$ is chosen as the index that minimizes the background contamination, defining the discrimination observable as $M_{1}$.
@@
The background contamination and signal efficiency can be computed by scanning over all possible values of the $M_{1}$ threshold, as exemplified in Fig.˜10. The contamination decreases with increasing primary energy, reflecting the increasing number of muons produced in hadronic air showers at higher energies. In addition, a larger signal efficiency can be reached at the expense of a larger background contamination.
@@
We note that the background contamination and signal efficiency shown in Fig.˜10 are obtained assuming a complete UMD hexagon, e.g., the north-west hexagon in Fig.˜1, left. We verified that the choice of $b=1$ remains effective even when considering missing UMD stations in the first ring of neighbors. However, the discrimination power, i.e. both the background contamination and signal efficiency, varies depending on the real configurations of the detector, as discussed in the next section.
@@ 5 Selection of photon events in data
In this section, we study the application of the discrimination method to data, considering the actual configuration of the detector. Given that the fraction of photon primaries expected to be present in the cosmic-ray flux may be well below $10^{-5}$, the standard event reconstruction, fine-tuned to hadronic events, is employed for the data. The event selection follows the procedure described in Section˜2. Both the estimated shower size and zenith angle are used to assign photon-equivalent energy to each event.
The observable $M_{1}$ quantifies the muon content of each event, as defined in Section˜4.2, and scales with the number of available UMD stations in the hottest hexagon. Air showers with low muon content, charactistic of photon-like events, would have smaller values of $M_{1}$ compared to hadronic events. However, if any UMD stations are missing, either due to not being deployed yet at the time of detection or to technical issues during data acquisition, less information on the muon content would be available. This results in smaller values of $M_{1}$ when fewer than seven UMD stations are operational, leading to an increase of the likelihood of misidentifying background events as photon-initiated. Thus, before applying the discrimination method to the selected data set, it is important to evaluate the impact of missing UMD modules across the seven stations of a hexagon.
Due to the irregular placement of UMD stations in the SD-433, as seen in Fig.˜1, left, the muon component arriving at the ground is sampled by a variable number of UMD stations. Therefore, the threshold value of $M_{1}$ to classify an event as a photon candidate, i.e., the photon candidate cut $M_{1}^{\text{cut}}$, is obtained according to the possible configurations in the hottest hexagon. The events are classified in six categories, as listed in Table˜4, depending on the following three criteria: the number of available UMD stations (i.e., stations with at least one active module), the total area spanned by these stations, and the relative position of any missing stations (i.e., stations without active modules). The hottest station must contain three active UMD modules regardless of the event category.
Cat.
No. of first-ring
UMD stations
Minimum UMD area
in first ring (m2)
Missing UMD
stations in first ring
No. of
events
I 66 $190$ 0 3,295
II 55 $140$ 11 1,491
III 44 $110$ 22 non-NN 8,417
IV 44 $110$ 22 NN 298
V 33 $80$ 33 non-NN 1,016
VI 33 $80$ 22 NN + $1$ non-NN 1,402
@@
The discrimination power of a hexagonal grid that mimics the SD-433 is evaluated through simulations (see Section˜3.1), assuming each SD station is paired with a 30 m2 UMD station, except for the aforementioned 50 m2 station. UMD stations or individual modules are randomly masked to simulate missing measurements and assess the discrimination power for each event category. $M_{1}^{\text{cut}}$ is defined as the threshold value at which the signal efficiency reaches $50\%$, determined as the median of the $M_{1}$ distribution for simulated photon events in energy bins. Such signal efficiency is suitable to reach background contamination levels smaller than $10^{-4}$ in all event categories. Therefore, an event is tagged as a photon candidate if $M_{1}<M_{1}^{\text{cut}}$. As shown in Fig.˜11, left, $M_{1}^{\text{cut}}$ increases with primary energy and with the number of UMD stations in the hottest hexagon . The evolution of $M_{1}^{\text{cut}}$ is described by a function of the logarithmic energy for each event category and reported in Appendix˜B. As depicted in Fig.˜11, right, the contamination remains below $10^{-5}$ for events in category I, while it increases as the number of first-ring UMD stations decreases, with the relative position of missing UMD stations having a significant impact. For instance, the contamination in events with two missing neighboring stations is two to four times larger than in events with non-neighboring missing stations. The background contamination is modeled using an exponential function of logarithmic energy, as detailed in Appendix˜B.
@@
Figure 11: The photon candidate cut (left) and the background contamination at $50\%$ signal efficiency (right) in terms of the photon-equivalent energy for each of the event categories cited in Table˜4. The dashed lines show examples of the fitted power-law and exponential models of the logarithmic energy, respectively.
@@
The parameterization of the background contamination for each event category can be employed to obtain a conservative estimation of the number of hadronic events in data that would be misidentified as photon candidates. For instance, the expected number of fake photon events in the category I subset can be computed as $(4.04\pm 0.04)\times 10^{-3}$ by integrating the blue curve in Fig.˜11, right, above $10^{16.7}$ eV [68]. Since 3,295 events above $10^{16.7}$ eV fall in that event category, this is equivalent to one fake photon event approximately every 815,650 events [^5] The expected number of fake photon events in the category IV and VI subsets, corresponding to the configurations with the largest background contamination, is $3.12\times 10^{-3}$ and $8.20\times 10^{-3}$ respectively. It can be translated to one fake photon event every 95,460 and 170,890 events in each of the subsets.. Considering that the category I events were acquired over $\sim 10\,\mathrm{months}$ of the observation time, this suggests that one fake photon event would be expected every 106 years per hexagon.
@@
Given the low probability of observing a background event consistent with the photon hypothesis, a fraction of the selected data, referred to as the burnt data sample, is unblinded to check that the $M_{1}$ distribution in data agrees with simulations, before carrying on the unblinding of the full selected data set. The burnt data represent a fraction $f_{\text{burnt}}=0.1$ of the selected data set. The $M_{1}$ distributions of burnt data for each event category are compared with the simulated distributions, as exemplified in Fig.˜12. Simulated events are weighted by $E_{\gamma,\text{eq}}^{-2}$ to obtain energy-integrated distributions, while no weighting is applied to the burnt data sample. As expected, the $M_{1}$ distribution for the burnt data is compatible with a hadronic origin. It is important to note that the photon-equivalent energy assigned to a data event is around $\sim 30\%$ higher than that assigned using the hadronic-optimized energy scale. This discrepancy arises from the larger muon content in cosmic-ray showers compared to the photon showers used to define the photon-equivalent energy calibration. As a result, the reference muon density $\rho_{\text{pr}}$ used in the $M_{1}$ definition (see Eq.˜4.1) is shifted to higher values in data compared to simulations, leading to smaller $M_{1}$ values in the former case. Thus, the distribution for data is artificially shifted towards the simulated proton distribution. Nevertheless, the $M_{1}$ observable is optimized to discriminate between events with significantly different muonic components -such as photons versus all hadronic species- due to the logarithm scaling in its definition, making it largely insensitive to the primary hadronic mass. This limitation does not impact the search for photon events in data, as discussed in the next section.